OpenStudy (anonymous):

Caitlin had $402 in her bank account. She withdrew $15 each week to pay for a swimming lesson. She now has $237. a. Write an equation that can be used to find the number of swimming lessons that she paid for. b. How many swimming lessons did she pay for? c. At the time she had $237, the cost of a lesson rose to $19. How many lessons can she pay for with her remaining $237?

OpenStudy (mertsj):

402-15w=237

OpenStudy (mertsj):

w=11

OpenStudy (experimentx):

let x be money on bank, and w be week, x = 402 - w*15 would be your equation. for b, put x = 237 and solve for w, c) change the equation to x = 237 - 19 w , put x = 0, and calculate w

OpenStudy (anonymous):

willyou please label what part of the question this is for there is a b c

OpenStudy (mertsj):

\[\frac{237}{19}=12\]

OpenStudy (mertsj):

You can figure that out.

OpenStudy (anonymous):

i dont know how to do what you are asking me to do wow i just confused myself writting this message

OpenStudy (experimentx):

a) x = 402 - w*15

OpenStudy (anonymous):

i already have that part just dont know how to do b and c

OpenStudy (mertsj):

To get b, just solve the equation from a

OpenStudy (mertsj):

If you have $50, and concert tickets are $10, how many concert tickets can you buy?

OpenStudy (anonymous):

thats where im confused i dont know how to solve it

OpenStudy (anonymous):

5 tickets

OpenStudy (mertsj):

How did you get that?

OpenStudy (anonymous):

50/10

OpenStudy (mertsj):

So if you have $237 and swimming lessons are $19 each, how many swimming lessons can you buy?

OpenStudy (anonymous):

12.47

OpenStudy (mertsj):

Well, I doubt anyone is going to give you .47 swimming lessons so the answer is that you could buy 12 lessons and you would have a little money left over.

OpenStudy (anonymous):

is this for c

OpenStudy (mertsj):

What is the question in c?

OpenStudy (anonymous):

At the time she had $237, the cost of a lesson rose to $19. How many lessons can she pay for with her remaining $237?

OpenStudy (anonymous):

if 12 is for c then what do i do for b

OpenStudy (mertsj):

So, how many lessons can she pay for with $237?

OpenStudy (anonymous):

12

OpenStudy (mertsj):

So you have just answered your own question.

OpenStudy (mertsj):

Solve the equation from a. That will be the answer to b.

OpenStudy (mertsj):